Differential Equations and Matrix Algebra problem:

Problem 1 Find the general solution 2(t) = ( 2) -(0) 0 1 2. rt) 0 1 0 3. x(t) = (* (1)«ce)

Can you please show how to do numbers 2 and 3? Could you show once you find the eigenvalues, the steps you take for the Gaussian Elimination and row reducing to get the eigenvectors? I'm having trouble with the Gaussian Elimination portion of the problem, trying to get the bottom row of the matrix to be all zeros. For problem 3, I found the eigenvectors when lambda is equal to 0, but I'm stuck with lambda equal to 2. Thanks.
Answer #1

Q Given that the 2017) 0 Delt) - - 3 2 2-2 x!(t) = A 210) characteristic Squation 1A-Il fo -3-1 r2 SO JI - 189] 22 -2-X . (-22 A-NI) 2) 2-1 R₂ 7 2 R₂ +VER, - (2) (23) au 10 ) .-28,+EX2=0 let take *2=k +22;= f 82% 21=22x 2 so Hitam); WEK (A) = x( Vd=-4 (A+UN) X244) = 0 Atux.= F3+4 & r. -2+4 ) Atuy R2 누 ( a 60%]K]=6 R₂ - V2RI dit (222=0 42 =k =) di = - VZH₂ x=-Vek - so;1-) - (-1) (02-1) -1 (0-1)+1(1+0=0 -03+titltltd.co 3 do +30+2=0 03-31-220 of deu tien 7+8=2=0 day 2 1o-3 -2 O-II | -1 -2 (2²-ar 212 23 (8) Orbitary Constant (f(n) {n} 3-1=2 KithfX₂=0 X2 = ki = kg X2= 2. = KitK2) = fkitka Xilt ci Kq d2=-2 then (A-II)21,30 X250 alt)= ( ; 있 23=0 heuce the general solution ip Xit) X lt) = & +60) et pkite tge 248 ki K2 x It) = (0,+ c t )e-t(K,1 © xlt) = 6; ;) 11 dIt) = A alt) IA= A=(-1) O Sigoe Vector Congesponding to 1-0,23 here (A-VI) & co, 4,50 ; (A-AI) = (-1) tk Wilt) = x=2 then (A-XI) = X/16) -- *11) (4-3) =<:: 603) 3] =6) R₂ > R2 i Ri -xit inco X =k ide x= X = ik hence thee xus] =>